OpenStudy (anonymous):

Gabriella went skiing. She paid $35 to rent skis and $15 an hour to ski. If she paid a total of $95, how many hours did she ski? Write an equation to represent the situation and solve it.

OpenStudy (anonymous):

@563blackghost

563blackghost (563blackghost):

She paid $35 to rent the skis but it costs 15 per hour to ski so our equation looks like.... \(\Huge{35+15h}\) Now she paid a total of $95 so it would then be... \(\Huge{35+15h=95}\) To solve we would subtract 35 from both sides.... \(\Huge{15h=60}\) in which you divide by 15 to get h by itself....

OpenStudy (anonymous):

What do you divide by 15? @563blackghost

563blackghost (563blackghost):

\(\Huge{\frac{15h}{15}=\frac{60}{15}}\)

OpenStudy (anonymous):

So, 60/15 will be the final answer? @563blackghost

OpenStudy (anonymous):

When you divide it @563blackghost

563blackghost (563blackghost):

Nope you have divide them both that would be your answer.... \(\Huge\color{red}{h=?}\)

OpenStudy (anonymous):

Thank You @563blackghost

563blackghost (563blackghost):

\(\Huge\cal\rlap{\color{red}{No~Problem!:)}}{\color{blue}{\;No~Problem!:)}}\)